# Radar-Guided Polynomial Fitting for Metric Depth Estimation

### v2

@@ Appendix G Additional Comparisons
DPT (Ranftl et al., 2021) 4117.9 5498.9
Depth Anything (Yang et al., 2024) 3270.5 4411.9
Depth Pro (Bochkovskii et al., 2024) 3275.9 5936.7
UniDepth (Piccinelli et al., 2025) 2605.6 5691.0
@@
POLAR w/ UniDepth 1500.1 3960.5
Leveraging Radar. Tab. 13 shows that replacing radar points with learnable, dataset-specific points worsens MAE by 28.0% and RMSE by 29.9%, demonstrating that we indeed leverage the radar inputs effectively. As additional evidence, Tab. 12 shows that our method, evaluated zero-shot cross-dataset, achieves comparable or better performance than baselines trained on the target datasets. Tab. 12 shows we are more robust to reduced radar point density at inference, i.e., less performance degradation than the baseline method RadarCam-Depth.
Table 11: Zero-shot generalization.
@@
Table 12: Reduced radar point density.
@@
Adapters. Tab. 13 shows that recent adapter-based finetuning methods LoRA (Hu et al., 2022) and ViT-Adapter (Chen et al., 2022), even with a post-hoc linear fit to projected radar points, do not outperform us.
Regression Baselines. Isotonic regression and monotone spline fitting methods are natural baselines. Tab. 13 shows these methods for regressing projected radar points on MDE predictions do not outperform us. We hypothesize that this is due to noise in radar points that can be mitigated by learning (Sec. 3.2).
Our learned polynomial fit may, in principle, introduce unwanted inversions of initially correct MDE predictions. POLAR successfully mitigates this effect through the proposed novel first-derivative regularization term (see Sec. 3.4 and Eqs. 67), which effectively constrains such inversions. To quantify this, we compute Kendall’s $\tau$ coefficient between predicted and ground-truth depths. Our method achieves the highest monotonicity ($\tau=0.969$) over regression baselines Isotonic Regression ($\tau=0.871$), PCHIP ($\tau=0.758$), and Cubic Hermite Spline ($\tau=0.736$). The raw MDE predictions do exhibit monotonicity with respect to ground truth ($\tau=0.957$), but our polynomial transformation increases it, indicating that we correct unwanted inversions. To assess statistical significance, we compute Kendall’s $\tau$ over 30 bootstrap samples for both our method and the raw MDE predictions. A two-sample t-test reveals a statistically significant difference in mean monotonicity ($p=0.012$).
@@ Appendix H Proof: Limitations of Global Scale and Shift
Consider the following three (MDE prediction $\hat{d}$, ground truth $d$) pairs from the nuScenes dataset, specifically from the image shown in Fig. 3:
@@
$7=5\alpha+\beta,\quad 9=10\alpha+\beta.$

### v3

@@ Appendix F Evaluation Metrics and Implementation
DPT [ranftl2021vision] 4117.9 5498.9
UniDepth [piccinelli2025unidepthv2] 2605.6 5691.0
Depth Anything [depth_anything_v2] 3270.5 4411.9
Depth Pro [depth-pro] 3275.9 5936.7
@@
Table 11: MDE backbone comparative studies on View-of-Delft.
@@
Table 12: Zero-shot generalization.
@@
Appendix G Additional Comparisons
Full MDE Comparisons. For VoD (see Tab. 11), POLAR w/ UniDepth achieves the best performance in MAE, improving over raw UniDepth predictions by 42.4% and over RadarCam-Depth w/ UniDepth by 32.7%, while POLAR w/ Depth Anything achieves the best performance in RMSE, improving over RadarCam-Depth w/ Depth Anything by 37.6% and over raw Depth Anything predictions (inverted and median scaled) by 10.4%.
Leveraging Radar. Tab. 14 shows that replacing radar points with learnable, dataset-specific points worsens MAE by 28.0% and RMSE by 29.9%, demonstrating that we indeed leverage the radar inputs effectively. As additional evidence, Tab. 12 shows that our method, evaluated zero-shot cross-dataset, achieves comparable or better performance than baselines trained on the target datasets.
Radar Sparsity. Tab. 13 shows we are more robust to reduced radar point density at inference, i.e., less performance degradation than the baseline method RadarCam-Depth.
Adapters. Tab. 14 shows that recent adapter-based finetuning methods LoRA [hu2022lora] and ViT-Adapter [chen2022vision], even with a post-hoc linear fit to projected radar points, do not outperform our method.
Regression Baselines. Isotonic regression and monotone spline fitting methods are natural baselines. Tab. 14 shows these methods for regressing projected radar points on MDE predictions do not outperform us. We hypothesize that this is due to noise in radar points that can be mitigated by learning (Sec. 3.2 from the main paper).
Our learned polynomial fit may, in principle, introduce unwanted inversions of initially correct MDE predictions. POLAR successfully mitigates this effect through the proposed novel first-derivative regularization term (see Sec. 3.4 and Eqs. 6, 7 from the main paper), which effectively constrains such inversions. To quantify this, we compute Kendall’s $\tau$ coefficient between predicted and ground-truth depths. Our method achieves the highest monotonicity ($\tau=0.969$) over regression baselines Isotonic Regression ($\tau=0.871$), PCHIP ($\tau=0.758$), and Cubic Hermite Spline ($\tau=0.736$). The raw MDE predictions do exhibit monotonicity with respect to ground truth ($\tau=0.957$), but our polynomial transformation increases it, indicating that we correct unwanted inversions. To assess statistical significance, we compute Kendall’s $\tau$ over 30 bootstrap samples for both our method and the raw MDE predictions. A two-sample t-test reveals a statistically significant difference in mean monotonicity ($p=0.012$).
@@ Appendix H Proof: Limitations of Global Scale and Shift
Consider the following three (MDE prediction $\hat{d}$, ground truth $d$) pairs from the nuScenes dataset, specifically from the image shown in Fig. 4 from the main paper:
@@
$7=5\alpha+\beta,\qquad 9=10\alpha+\beta.$